# A Cascaded Random Access Quantum Memory

### v1

@@ IV Random Access Quantum Memory Performance
Figure 3: Random access quantum memory operation. (a) RAM RB circuit on $7$ storage modes. Each storage access gate is highlighted in a different color, containing a storage-buffer swap (read), a buffer gate $G_{i}^{n}$ (address), and a buffer-storage swap (write). The storage access gate in the initialization stage does not contain the storage-buffer swap (read), as the storage modes are initialized in the ground state. At the end of the circuit, the target storage mode information is swapped into the buffer for readout. (b) Measured random read fidelity for $7$ storage modes. The reference RB is the buffer mode’s RB in the ${\left|0\right\rangle,\left|1\right\rangle}$ subspace. (c) Measured average random read infidelity at different RAQM sizes. The region below the depolarization error correction [29, 30, 31, 32] threshold for surface code is highlighted.
After demonstrating single-mode memory performance, we extend our analysis to RAQM performance. An ideal RAQM should have high-fidelity random read/write operations with minimal disturbance to other stored states. In our system, the random read/write is the buffer-storage swap. To efficiently evaluate RAQM performance, we perform multiplexed RAM RB by applying memory gates cyclically across the storage modes (See Fig. 3(a)). We vary our RAQM size and show that our memory is below the memory error correction threshold [29, 30, 31, 32].
@@
We first perform the multiplexed RAM RB to measure the random read/write fidelity. Fig. 3(a) shows the corresponding circuits. We sequentially perform RB gates on each storage mode. Each storage mode tracks its respective RB gate sequence during this multiplexing process. Each storage gate is highlighted in a box, which contains a buffer-storage swap (read), a buffer gate $G_{j}^{n}$ (address), and a storage-buffer swap (write). For smaller RAQM sizes, RB gates are applied cyclically to the corresponding subsystem.
@@
The measured random read fidelity for a size-7 RAQM is shown in Fig. 3(b). For each mode $S_{i}$, the depolarization parameter $P_{S_{i}}$ is extracted by fitting the process fidelity after different RAQM cycles. For a given mode $S_{i}$, a single cycle in a size-7 memory consists of 14 gates: two gates (read and write) are performed on $S_{i}$, while the other 12 gates are performed on the other storage modes. The latter are effectively identity gates on $S_{i}$. However, due to decoherence and modes’ cross-Kerrs, these identity gates have nonzero infidelities. $S_{i}$’s random read fidelity $\bar{F}_{BS_{i}}$ is
Here, we isolate the buffer mode’s control fidelity, with its depolarization parameter defined as $P_{B}$. $P_{B}$ is measured independently using single-mode RB on the buffer, involving only $G_{i}^{n}$. The factor $2/3$ comes from the fact that $3$ levels are used in the RAQM controlling transmon [33]. The measured $S_{i}$’s random read fidelity $\bar{F}_{BS_{i}}$ ranges between $(98.74\%\pm 0.04\%)$ and $(98.89\%\pm 0.05\%)$. Overall, any read/write operations on a size-7 RAQM imparts an error $\sim 1.2\%$ on each mode.
Unlike its classical counterpart, RAQM can have an additional error channel from many-body interactions. To explore their effects on memory scalability, we extracted the random read infidelity at different RAQM sizes (See Fig. 3(c)). In our device, a size-7 memory is the largest RAQM whose infidelity per mode after a cycle –consisting of a read and a write on each mode– is ($1-0.9874^{14}=16.27\%$) below the depolarization error threshold for the surface code ($\sim 17\%$[29, 30, 31, 32]. Larger RAQM requires more operations to complete one cycle, resulting in the power-law decay of the depolarization error threshold. As the memory size increases, the random read infidelity is dominated by idle errors, indicating that lower idle errors are key to RAQM’s scalability.
@@
Figure 4: RAQM crosstalk benchmarking. (a) Measured cross-Kerr values between storage and buffer modes (log-scale). (b) Three RAQM periods for a target storage mode $S_{i}$. Buffer gates $G_{1}$ and $G_{2}$ are applied during $S_{i}$’s active and inactive access periods. (c) $BS_{i}$ active access error due to states stored in $S_{j}$. The error rate is extracted using the bottom RB sequence. (d) The additional dephasing rate during $S_{j}$’s inactive access period, where the state stored in spectator modes $S_{i}$ $(i\neq j)$ undergoes random gates. This acts as a dephasing channel to $S_{j}$, with the rate measured using the bottom sequence. (e) The many-body dephasing rate during the idle period, where the state stored in $S_{i}$ experiences additional dephasing due to cross-Kerr interactions between storage mode. This rate is measured using the bottom sequence. (f) The random read fidelity error budget. The dominant error channels come from storage decay and $BS_{i}$ cross-Kerr interactions. Details of the error budget are shown in Supplementary Section I.
@@
V Many-body Interactions in Random Access Quantum Memory
While bosonic modes should be transparent to each other, nonlinear control elements in RAQM induce cross-Kerr interactions between modes, which can lead to many-body interactions. These interactions effectively dephase the stored states, even for ideal cavities. We characterize the strength of these interactions in Fig. 4(a).
The cross-Kerr interactions between modes lead to errors of different flavors depending on RAQM operation. To characterize these errors, we divide RAQM control into three periods (see Fig. 4(b)) referenced to a given mode $S_{i}$: (i) Active access period, occurs when $S_{i}$ is accessed; (ii) Inactive access period, occurs when other modes $S_{j}(j\neq i)$ are accessed; (iii) Idle period, occurs when none of the modes are accessed (i.e. during transmon operations). Active errors occur twice per mode (read and write) per cycle, whereas inactive access and idle errors scale with the memory size.
@@ V Many-body Interactions in Random Access Quantum Memory
During $S_{i}$’s active access period, the dominant crosstalk is the state-dependent access error (see Fig. 4(c)). Populations in other storage modes shift the $BS_{i}$ swap frequency by buffer-storage cross-Kerrs, reducing $S_{i}$’s access fidelity. To benchmark this error per active swap, we perform RB on $BS_{i}$ swap gates with different states stored in the other modes. Details about the RB fitting are shown in Supplementary Section G. Overall, the $BS_{i}$ gate incurs an additional error of $1.069\%$ due to populations in spectator modes.
@@
During the idle period, the main crosstalk on $S_{i}$ is the storage many-body dephasing (see Fig. 4(e)). When all storage modes are occupied with unknown states, the storage-storage cross-Kerrs are converted into a many-body dephasing channel on $S_{i}$. To extract this error, we perform the Ramsey experiment on $S_{i}$ while storing identical cardinal states of the Bloch sphere in the other modes (See Supplementary Section G for analysis). This storage many-body dephasing increases the storage Ramsey dephasing rate by $0.223\,\mathrm{kHz}$ on average.
The modes’ decoherence and many-body interactions are the main error channels for memory operations. Fig. 4(f) shows the error budget of random read fidelity $\bar{F}_{BS_{i}}$ per gate of the size-7 RAQM. The state-dependent access error is determined by combining its swap infidelity from Sec. III with the state-dependent access error from its spectator modes. The decay, dephasing, and spectator-access dephasing errors for $S_{i}$ are computed throughout its inactive access period. The many-body dephasing error is computed over the full duration of an RAQM cycle. Although the spectator-access dephasing error is dominant, it is due to the long inactive access period compared to other stages of RAQM control. Details about the error budget are shown in Supplementary Section I.
Although the coupler and transmon have typical anharmonicities, the additional buffer layer shields the propagation of Kerrs, rendering the modes nearly transparent to each other. While many-body interactions are still the dominant error source, they are comparable to modes’ decoherence [19] and do not substantially impact the scalability of random read fidelity with memory size, as shown in Fig. 3(c). Thus, the cascaded architecture improves the scalability of RAQM.
@@ VI Conclusion and Outlook
While the fidelity demonstrated here is already above the nominal threshold for fault-tolerant memory, further improvement would enable practical use. The current dominant error, many-body interactions between buffer and storage modes, can be suppressed significantly using a more linearized coupler design with several junctions in series [34] and/or a balanced design [35]. Niobium cavities [11, 36, 37] have been proven to have $T_{1}$ beyond $20$ ms, which will significantly reduce the intrinsic decoherence error of the memories. Compact on-chip storage options using nanomechanical resonators [22] and spin memories [23] can also provide longer coherence. With these improvements, the random read fidelity across all storage modes can exceed $99.9\%$ and support over $100$ memory modes ($100$ times more logical qubits) compatible with the error correction, which can be randomly accessed and operated on with transversal two-qubit gates. The architecture is also compatible with hardware-efficient bosonic error correction codes [24] and biased-noise/dual rail encodings [25, 26] that can further improve the amount of memory storage and overhead of quantum error correction.
@@ Appendix A Device Parameters
$\Phi_{\text{dc}}=0.269\pi$ Symbol Frequency/$2\pi$ (GHz) $T_{1}\left(\mu\text{s}\right)$ $T_{R}\left(\mu\text{s}\right)$ $T_{\rm echo}\left(\mu\text{s}\right)$
@@ A.1 Coupler Dispersion
We will first discuss the dispersion of the coupler and the buffer mode. Fig. 5(a) shows the coupler and buffer mode frequencies at different DC flux points, $\Phi_{\text{DC}}$. For each value of $\Phi_{\text{DC}}$, the buffer mode is first initialized to $\left|1\right\rangle$ using the transmon-buffer modes $\left|f0\right\rangle\leftrightarrow\left|g1\right\rangle$ sideband. The RF flux modulation is then turned on for $1\,\mu\text{s}$, followed by another $\left|f0\right\rangle\leftrightarrow\left|g1\right\rangle$ sideband that maps the buffer state into the transmon $\{\left|g\right\rangle,\left|f\right\rangle\}$ subspace. This state is subsequently mapped to the $\{\left|g\right\rangle,\left|e\right\rangle\}$ subspace using a transmon $\pi$ pulse between $\left|f\right\rangle$ and $\left|e\right\rangle$. Blue regions in the plot indicate possible transitions between the buffer and other modes in the system. One of the most tunable traces corresponds to the coupler mode, which matches the black-box quantization results [38].
The $T_{1}$ data for the coupler and buffer modes is shown in Fig. 5(b). To measure buffer mode’s $T_{1}$, we prepare $|1\rangle$ in the buffer mode using the $|f0\rangle\leftrightarrow|g1\rangle$ sideband with the transmon, wait for a varied time, then apply the $|f0\rangle\leftrightarrow|g1\rangle$ sideband again and readout the qubit state. Since the coupler mode lacks a dedicated dispersive readout, we use a cascaded readout scheme: prepare $|1\rangle$ in $B$, apply an RF flux-modulated sideband between $|B\rangle$ and the coupler, transfer the excitation to the coupler, wait for a varied time, reverse the mapping, and finally read out the qubit state.
@@ A.2 Device Coherences
All modes’ frequencies and coherences at the operation point $\Phi_{\text{DC}}=0.269\pi$ (dashed line in Fig. 5(b)) are shown in Table 1. In our experiments, the $T_{1}$ of the storage modes is limited by the external drive pin length, resulting in an over-coupled regime where external coupling $Q_{c}$ is smaller than the modes’ intrinsic $Q_{i}$. We reduced the drive pin length in a subsequent cooldown to increase $Q_{c}$. In this new cooldown, we experimentally observed that $6$ out of the $7$ storage modes have $T_{1}$ exceeding $1\,\mathrm{ms}$. This coherence time is comparable to a bare Aluminum flute cavity [19]. Further improvements would require different cavity materials, such as niobium [37]. Detailed coherence time in the new cooldown is shown in Table 2.
Storage $1$ Storage $2$ Storage $3$ Storage $4$ Storage $5$ Storage $6$ Storage $7$
$T_{1}$ (ms) $1.153\pm 0.020$ $1.165\pm 0.019$ $1.138\pm 0.016$ $1.180\pm 0.009$ $1.086\pm 0.016$ $1.019\pm 0.014$ $0.629\pm 0.007$
@@ A.3 Self-Kerrs and Cross-Kerrs
Table 3 shows the cross-Kerrs and self-Kerrs measured in the experiments. The self-Kerrs $K$ of the two buffer modes $B_{i}$’s (including the unused $B_{2}$ mode in the experiments) are measured through two coherent cavity displacements gapped by a varied waiting time $\tau$ [19]. The displacement amplitude $\alpha$ is also varied. After each sequence, the population of $B_{i}$ in the state $\left|0\right\rangle$ is measured, and the oscillation frequency, $f_{d}$, is extracted for each $\alpha$. Under small-angle approximation $K|\alpha|^{2}t\ll\pi$  [40], the self-Kerr coefficient is determined as the fitted slope of $f_{d}$ to $\alpha$. $S_{i}$’s self-Kerrs are measured similarly, with the displaced state prepared in $B$ first and then swapped into $S_{i}$ through the $BS_{i}$ sideband. Cross-Kerrs between cavity modes are measured by performing Ramsey experiment on the target mode $S_{i}$ with the spectator mode $S_{j}$ initialized at $\left|0\right\rangle$ or $\left|1\right\rangle$. We measure the cross-Kerr strength between the coupler and $B_{i}(S_{5})$ by performing cavity mode spectroscopy while the coupler is at $\left|g\right\rangle$ or $\left|e\right\rangle$. The cross-Kerrs between the coupler and other storage modes are measured through the $BS_{i}$ sideband frequency difference when the coupler is at $\left|g\right\rangle$ or $\left|e\right\rangle$.

### v2

@@ IV Benchmarking Random Access Quantum Memory
Figure 3: Random access quantum memory operation. (a) RAM RB circuit on 77 storage modes. Each storage access gate is highlighted in a different color, containing a storage-buffer swap (read), a buffer gate $G_{i}^{n}$, and a buffer-storage swap (write). The storage access gate in the initialization stage does not contain the storage-buffer swap (read), as the storage modes are initialized in the ground state. At the end of the circuit, the target storage mode information is swapped into the buffer for readout. (b) Measured random read fidelity for $7$ storage modes. The reference RB is the buffer mode’s RB in the $\left|0\right\rangle,\left|1\right\rangle$} subspace. (c) Measured average random read infidelity at different RAQM sizes. The region below the depolarization error correction [66, 5, 13, 34] threshold for surface code is highlighted.
Having characterized the individual memory modes, we now evaluate the system’s performance as a random-access quantum memory (RAQM). An ideal RAQM enables high-fidelity read and write operations—implemented in our device as controlled buffer–storage swaps—on any selected mode while minimally disturbing all others. We benchmark these memory access operations across the full RAQM and discuss their implications for scalability.
@@
To assess RAQM performance, we perform multiplexed RB across all memory modes, as shown in Fig. 3(a). The benchmarking sequence consists of the cyclic application of memory access gates, where full memory access consists of a buffer-storage swap (read), a buffer gate $G_{j}^{n}$, and a storage-buffer swap (write). Each storage mode tracks its respective RB gate sequence during this multiplexing process. For smaller RAQM sizes, RB gates are applied cyclically to the corresponding subsystem.
@@
For a size-7 RAQM, one memory cycle consists of 14 gates: two (read and write) applied on a selected storage mode $S_{i}$, and 12 applied to other storage modes. Ideally, the latter are effectively identity gates on $S_{i}$. However, due to decoherence and modes’ cross-Kerrs, these identity gates have nonzero infidelities. From the RAQM RB data, we derive the random read fidelity $\bar{F}_{BS_{i}}$, which quantifies the average error per gate (Supplementary Section IX). As shown in Fig. 3(b), the measured fidelity $\bar{F}_{BS_{i}}$ for $S_{i}$ ranges between $(98.74\%\pm 0.04\%)$ and $(98.89\%\pm 0.05\%)$. Overall, any read/write operation on a size-7 RAQM imparts an error $\sim 1.2\%$ on each mode.
These access errors necessitate periodic refresh of the memory using quantum error correction. We therefore benchmark the RAQM in QEC-relevant terms. Specifically, we identify the largest memory size for which errors accumulated over a full memory cycle remain below the QEC threshold. We address this by extracting the random read infidelity for different RAQM sizes (See Fig. 3(c)). In our device, a size-7 memory is the largest RAQM that remains below the surface-code depolarization threshold of about $17\%$ [66, 5, 13, 34]. After a full cycle, which consists of a read and a write on each storage mode, the infidelity per mode is $1-0.9874^{14}=16.27\%$.
For purposes of benchmarking the memory, the threshold above assumes ideal processor gates and stabilizer measurements, with errors arising only from memory read/write operations. Notably, this threshold is higher than that for two-qubit gates, as memory swaps are performed only twice per QEC round. Larger RAQM requires more operations to complete one cycle, resulting in the power-law decay of the depolarization error threshold (See Fig. 3(c)).
These experiments reveal insights into the scalability of the RAQM architecture shown in Fig. 1(a). An RAQM substantially reduces the number of control lines only when the coherence of the memory exceeds that of the processor and the processor–memory swap error remains below the threshold. Increasing the number of memory modes enhances the logical qubit capacity without requiring additional control lines, but also amplifies idle errors. Beyond a certain memory size, the resources required to correct these additional errors can outweigh the control-line savings. Thus, suppressing idle errors is crucial for RAQM scalability. The optimal memory size is discussed further in Supplementary Section XII.
@@
Figure 4: RAQM crosstalk benchmarking. (a) Measured cross-Kerr values between storage and buffer modes (log-scale). (b) Three RAQM periods for a target storage mode $S_{i}$. Buffer gates $G_{1}$ and $G_{2}$ are applied during $S_{i}$’s active and inactive access periods. (c) $BS_{i}$ active access error due to states stored in $S_{j}$. The error rate is extracted using the bottom RB sequence. (d) The additional dephasing rate during $S_{j}$’s inactive access period, where the state stored in spectator modes $S_{i}$ $(i\neq j)$ undergoes random gates. This acts as a dephasing channel to $S_{j}$, with the rate measured using the bottom sequence. (e) The many-body dephasing rate during the idle period, where the state stored in $S_{i}$ experiences additional dephasing due to cross-Kerr interactions between storage mode. This rate is measured using the bottom sequence. (f) The random read fidelity error budget. The dominant error channels come from storage decay and $BS_{i}$ cross-Kerr interactions. Details of the error budget are shown in Supplementary Section X.
@@
V Exploring Many-body Interactions in Random Access Quantum Memory
While bosonic modes in a linear cavity are naturally non-interacting, the introduction of nonlinear control elements induces parasitic cross-Kerr interactions between all modes (Fig. 4(a)). These interactions lead to state-dependent frequency shifts and many-body dephasing. To characterize the impact of these error channels, we divide the RAQM operation cycle into three distinct regimes with respect to a target storage mode $S_{i}$ (Fig. 4(b)). First, the Active Access period occurs when $S_{i}$ is actively swapped to or from the buffer. Second, the Spectator Access period covers the duration when another mode $S_{j\neq i}$ is being swapped, leaving $S_{i}$ idle but exposed to buffer activity. Finally, the Idle period corresponds to times when the buffer is empty, such as during processor computation. Crucially, Active errors occur a fixed number of times (twice) per cycle, while Spectator and Idle errors accumulate linearly with the memory size $N$.
@@ V Exploring Many-body Interactions in Random Access Quantum Memory
During $S_{i}$’s active access period, the dominant crosstalk is the state-dependent access error (see Fig. 4(c)). Populations in other storage modes shift the $BS_{i}$ swap frequency by buffer-storage cross-Kerrs, reducing $S_{i}$’s access fidelity. To benchmark this error per active swap, we perform RB on $BS_{i}$ swap gates with different states stored in the other modes. Details about the RB fitting are shown in Supplementary Section VII. Overall, the $BS_{i}$ gate incurs an additional error of $1.069\%$ due to populations in spectator modes.
@@
During the idle period, the main crosstalk on $S_{i}$ is the storage many-body dephasing (see Fig. 4(e)). When all storage modes are occupied with unknown states, the storage-storage cross-Kerrs are converted into a many-body dephasing channel on $S_{i}$. To extract this error, we perform the Ramsey experiment on $S_{i}$ while storing identical cardinal states of the Bloch sphere in the other modes (See Supplementary Section VII for analysis). This storage many-body dephasing increases the storage Ramsey dephasing rate by $0.223\,\mathrm{kHz}$ on average.
The modes’ decoherence and many-body interactions are the main error channels for memory operations. Fig. 4(f) shows the error budget of random read fidelity $\bar{F}_{BS_{i}}$ per gate of the size-7 RAQM. The state-dependent access error is determined by combining its swap infidelity from Section III with the state-dependent access error from its spectator modes. The decay, dephasing, and spectator-access dephasing errors for $S_{i}$ are computed throughout its inactive access period. The many-body dephasing error is computed over the full duration of an RAQM cycle. Although the spectator-access dephasing error is dominant, it is due to the long inactive access period compared to other stages of RAQM control. Details about the error budget are shown in Supplementary Section X.
[…]
@@ VI Conclusion and Outlook
While the fidelity demonstrated here is already above the nominal threshold for fault-tolerant memory, further improvements can enable practical use. The current dominant error, many-body interactions between buffer and storage modes, can be suppressed significantly using a more linearized coupler design with several junctions in series [69] and/or a balanced design [10, 41]. Niobium cavities [53, 45, 49] have been proven to have $T_{1}$ beyond $20$ ms, which will significantly reduce the intrinsic decoherence error of the memories. Compact on-chip storage options using nanomechanical resonators [14] and spin memories [52] can also provide longer coherence. With these improvements, the random read fidelity across all storage modes can exceed $99.9\%$ and support over $100$ memory modes ($100$ times more logical qubits) compatible with the error correction, which can be randomly accessed and operated on with transversal two-qubit gates. The architecture is also compatible with hardware-efficient bosonic error correction codes [48, 44, 7, 56, 46] and biased-noise/dual rail encodings [59, 35, 12, 32] that can further improve the amount of memory storage and overhead of quantum error correction.
@@ Appendix A Device Parameters
[…]
@@ A.1 Coupler Dispersion
We will first discuss the dispersion of the coupler and the buffer mode. Supplementary Figure 5(a) shows the coupler and buffer mode frequencies at different DC flux points, $\Phi_{\text{DC}}$. For each value of $\Phi_{\text{DC}}$, the buffer mode is first initialized to $\left|1\right\rangle$ using the transmon-buffer modes $\left|f0\right\rangle\leftrightarrow\left|g1\right\rangle$ sideband. The RF flux modulation is then turned on for $1\,\mu\text{s}$, followed by another $\left|f0\right\rangle\leftrightarrow\left|g1\right\rangle$ sideband that maps the buffer state into the transmon $\{\left|g\right\rangle,\left|f\right\rangle\}$ subspace. This state is subsequently mapped to the $\{\left|g\right\rangle,\left|e\right\rangle\}$ subspace using a transmon $\pi$ pulse between $\left|f\right\rangle$ and $\left|e\right\rangle$. Blue regions in the plot indicate possible transitions between the buffer and other modes in the system. One of the most tunable traces corresponds to the coupler mode, which matches the black-box quantization results [47].
The $T_{1}$ data for the coupler and buffer modes is shown in Supplementary Figure 5(b). To measure buffer mode’s $T_{1}$, we prepare $|1\rangle$ in the buffer mode using the $|f0\rangle\leftrightarrow|g1\rangle$ sideband with the transmon, wait for a varied time, then apply the $|f0\rangle\leftrightarrow|g1\rangle$ sideband again and readout the qubit state. Since the coupler mode lacks a dedicated dispersive readout, we use a cascaded readout scheme: prepare $|1\rangle$ in $B$, apply an RF flux-modulated sideband between $|B\rangle$ and the coupler, transfer the excitation to the coupler, wait for a varied time, reverse the mapping, and finally read out the qubit state.
@@ A.2 Device Coherences
All modes’ frequencies and coherences at the operation point $\Phi_{\text{DC}}=0.269\pi$ (dashed line in Supplementary Figure 5(b)) are shown in Supplementary Table 1. In our experiments, the $T_{1}$ of the storage modes is limited by the external drive pin length, resulting in an over-coupled regime where external coupling $Q_{c}$ is smaller than the modes’ intrinsic $Q_{i}$. We reduced the drive pin length in a subsequent cooldown to increase $Q_{c}$. In this new cooldown, we experimentally observed that $6$ out of the $7$ storage modes have $T_{1}$ exceeding $1\,\mathrm{ms}$. This coherence time is comparable to a bare Aluminum flute cavity [8]. Further improvements would require different cavity materials, such as niobium [49]. Detailed coherence time in the new cooldown is shown in Supplementary Table 2.
Storage 11 Storage 22 Storage 33 Storage 44 Storage 55 Storage 66 Storage 77
$T_{1}$ (ms) $153\pm 0.020 1.1$$65\pm 0.019 1.1$$38\pm 0.016 1.1$$180\pm 0.009$ $1.086\pm 0.016$ $1.019\pm 0.014$ $0.629\pm 0.007$
@@ A.3 Self-Kerrs and Cross-Kerrs
Supplementary Table 3 shows the cross-Kerrs and self-Kerrs measured in the experiments. The self-Kerrs $K$ of the two buffer modes $B_{i}$’s (including the unused $B_{2}$ mode in the experiments) are measured through two coherent cavity displacements gapped by a varied waiting time $\tau$ [8]. The displacement amplitude $\alpha$ is also varied. After each sequence, the population of $B_{i}$ in the state $\left|0\right\rangle$ is measured, and the oscillation frequency, $f_{d}$, is extracted for each $\alpha$. Under small-angle approximation $K|\alpha|^{2}t\ll\pi$  [11], the self-Kerr coefficient is determined as the fitted slope of $f_{d}$ to $\alpha$. $S_{i}$’s self-Kerrs are measured similarly, with the displaced state prepared in $B$ first and then swapped into $S_{i}$ through the $BS_{i}$ sideband. Cross-Kerrs between cavity modes are measured by performing Ramsey experiment on the target mode $S_{i}$ with the spectator mode $S_{j}$ initialized at $\left|0\right\rangle$ or $\left|1\right\rangle$. We measure the cross-Kerr strength between the coupler and $B_{i}(S_{5})$ by performing cavity mode spectroscopy while the coupler is at $\left|g\right\rangle$ or $\left|e\right\rangle$. The cross-Kerrs between the coupler and other storage modes are measured through the $BS_{i}$ sideband frequency difference when the coupler is at $\left|g\right\rangle$ or $\left|e\right\rangle$.